# citations.bib

### v1

@@ Appendix F Methodology
In this section, we describe our estimators algorithmically. The unified theoretical framework can directly be implemented into an algorithmic framework, which we depict in Fig. 5. The general procedure to generate the Shapley values in our framework requires three choices: (1) a sampling distributions on the index-sizes, (2) a strategy for sampling (with replacement, without replacement) and (3) an approximation method (least squares or matrix-vector). Least squares and matrix-vector estimation are reported in Section 2. The missing detail is how the sampling procedure is implemented (this is the middle column in Fig. 5). We report this in Algorithm 1 for with replacement sampling and in Algorithm 2.
Figure 5: The unified framework for estimating Shapley values with the proposed class of estimators. First, we define a distribution to apply to each bucket (i.e., to the selection of the bit vector to select - $p_{i}$ is the probability of sampling an item from bucket/coalition of size (or bit vector with Hamming weight) $i\in[d]$. Then we select a sampling strategy (with or without replacement). Finally, we select the approximation strategy. If we limit ourselves to $\ell_{2}$ squared and modified, and kernel distribution, this provides a total of $3\times 2\times 2=12$ estimators.
@@ F.1 With Replacement Estimators
Sampling with replacement to generate the sketch is a computationally efficient procedure that performs well in practice. However, if the number of samples $m>2^{d}$, the estimator will fail to compute exact Shapley values in general. We report the sampling procedure as implemented in our experimental evaluations in Algorithm 1.
@@
1:$d$: number of features $m$: number of samples expected by user, $p$ choice of distribution on $d$ buckets, maxval maximum value before Poisson approximation is applied (our algorithm uses $10^{10})$
2:$\bm{S}\bm{Z}\in\mathbb{R}^{m\times d},\bm{\tilde{W}}\in\mathbb{R}^{m\times m}$: sub-sampled $\bm{Z}$ matrix and weights
3:First sample with replacement $\lfloor\frac{m}{2}\rfloor$ from indices $i=1,...,d$, $j\in[m]$ each with probability $p_{1},...,p_{d}$; store the counts in $m_{i}$.
4:For each $i\in[\frac{d}{2}]$, construct $\bm{b}_{j}\in\{0,1\}^{d}$ by sampling with replacement $m_{i}$ bitstrings of size $i$, which is equivalent to sampling without replacement $i$ indices from $[d]$, uniformly; and generate, the complement $\bar{\bm{b}}_{i,j}=1-\bm{b}_{i,j}$.
5:Construct $\bm{w}_{j}=\left((d-1)2p_{j}\right)^{-1},$ so that:
@@ F.2 Without Replacement Estimators
At first glance, sampling based on the without replacement sampling scheme described in Section 2.3 (see also Section A.5) requires flipping $22^{[d]}-$ 2 coins. However, for the sampling distributions described in Section 2.2, given the size of a subset, the probability of picking any given subset of that size is constant. This observation is used in [musco2025provably] to avoid flipping exponentially many coins. In particular, one can determine which coins are heads by first determining how many heads there will be of a given subset size, and then picking the resulting subsets of this size uniformly at random. We describe a variant of [musco2025provably, Algorithm 2] in Algorithm 2.
Algorithm 2 Sampling without Replacement (paired sampling, modified from [musco2025provably])
1:$d$: number of features $m$: number of samples expected by user, $p$ choice of distribution on $d$ buckets, maxval maximum value before Poisson approximation is applied (our algorithm uses $10^{10})$
2:$\bm{S}\bm{Z}\in\mathbb{R}^{m\times d},\bm{\tilde{W}}\in\mathbb{R}^{m\times m}$: sub-sampled $\bm{Z}$ matrix and weights
3:Choose $\alpha$ such that
$\lfloor\frac{m}{2}\rfloor=\sum_{i=1}^{\lfloor\frac{d}{2}\rfloor}\min\left({d\choose i},\alpha 2p_{i}\right)$
@@
4:For (${d\choose i}<\texttt{maxval}$
$m_{i}\leftarrow\text{Binomial}\left({d\choose i},\min(1,2\alpha p_{i})\right)$
5:For (${d\choose i}\geq\texttt{maxval}$:
$m_{i}\leftarrow\text{Poisson}(2\alpha p_{i}),\forall i\in\left[\lfloor\frac{d}{2}\rfloor\right],$
and let $\tilde{m}=\sum_{i=1}^{\lfloor\frac{d}{2}\rfloor}m_{i}$.
6:Construct $\bm{b}_{j}\in\{0,1\}^{d}$, bitstring array of size $i\in[\frac{d}{2}],j\in[m_{i}]$ without replacement (e.g. using Fisher Yates shuffling or Algorithms 2,3 in [musco2025provably]) and generate the complement $\bar{\bm{b}}_{i,j}=1-\bm{b}_{i,j}$.
7:Construct $\bm{w}_{j}=\left((d-1)\min({d\choose j},2\alpha p_{j})\right)^{-1},$  [^9] In practice, for large (${d\choose j}$ this is just assumed to be $2\alpha p_{j}$. so that
@@ Appendix G Experimental Details
We use publicly available datasets for reproducibility; choosing particularly those available through the shap for their popularity, ease of use and for a direct comparison with [musco2025provably].
@@ G.1 Training Details
In this subsection, we detail the experimental design choices and hyperparameter for low Section 3.1 and Section 3.2, including implementation details, to promote reprodicibility.
@@ G.1.1 Low Dimensional Experiments
We refer to low dimensional experiments to the content of Section 3.1. For each dataset we train a decision tree from the xgboost. Specifically, we use the XGBRegressor class with 100 estimators and maximum depth of 10 for each task. We replacing missing values with the mean for that feature in the dataset. Note that the goal is not to achieve competitive performance but to rapidly train a model where the Shapley values can be computed exactly and efficiently. The train test splits are ordered 80/20 splits for all datasets; we pick as query and baseline points the first data points of the test and train datasets respectively.
@@ G.1.2 High Dimensional Experiments
We refer to high dimensional experiments to the content of Section 3.2. For the two classification tasks, we train a RandomForestClassifier from the sci-kit learn library. The random forest has maximum depth 15 and random state 42 for both tasks (MNIST and CIFAR-10). For both datasets, we pick as query and baseline points the first data points of the test and train datasets respectively.
For MNIST we use train test splits (80/20) with random state 42 using the train_test_split method on sci-kit learn. We achieve a test accuracy of $97.2\%$.
For the CIFAR-10 dataset, we use standard train test split from the torchvision library. We preprocess the input data by normalising. We achieve low accuracy $46.12\%$, as should be expected using basic tree-based models for CIFAR-10; however, using said models enables the exact computation of Shapley values, which we consider a more important aspect for the paper.
@@ G.2 Datasets
Adult. Demographic information about individuals collected from the 1994 U.S. Census database. It is used to predict whether a person earns more than \$$\$50,000$ dollars per year based on individual attributes: age, work, class, education, etc.,.
@@ G.2.2 Image Datasets
MNIST The MNIST dataset is a collection of handwritten digets (0-9); the classic task is to classify into their respective value. This is an incredibly popular dataset used for training and testing image classification algorithms, broadly available in several libraries (including sci-kit learn).
CIFAR-10 CIFAR-10 dataset is a classification task dataset where a collection of 60,000 images, with $32\times 32\times 3(=3072)$ dimensions, are mapped to target classes: cars, airplanes, birds, cats, deer, dogs, frogs, horses, ships and trucks. The standard split for this dataset is 10,000 test images and 50,000 training images, which we use in our experiments. This is considered a relatively challenging dataset for boosted trees and feedforward neural networks; good performance is achieved, however, for convolutional neural networks.
@@ Appendix H Extended Experimental Results
The goal of this section is to report the numerical results from our experiments. We first report the extended experiments from Section 3.1, followed by experiments in Section 3.2. Importantly, we share tables and plots containing our results.
@@ H.1 Low Dimensional Experiments
In low dimensional experiments, as described in Appendix G and Section 3.1, we compute the mean squared error (normalized by norm of exact Shapley values) as the median of 100 random seeds (0-99). We also report the average and interquartile ranges for each of the experiments. These results are summarized in Table 2 (median), Table 3 (lower quantile), Table 4 (upper quantile), and Table 5 ($b_{0}$ values). In each of those tables we report the values for selected number of samples: for IRIS we report $m=10$; Adult, California and Diabetes we show results for $m=64$; for Communities, Correlated, Independent, and NHANES we report $m=16384$. Reporting higher $m$ values in the low dimensional datasets would result in comparing numerical errors of exact estimators, whereas we want to assess the estimators ability to approximate the true values.
@@ H.2 High Dimensional Experiments
For high dimensional experiments, as described in Appendix G and Section 3.2, we compute the the mean squared error (normalized by norm of exact Shapley values) as the average of 10 test points from the respective datasets. We also report quantiles and median in Table 6.
@@ H.3 Faithfulness Experiments
For each experiment in the high dimensional setting, we compute the insertion and deletion curves as reported in Table 7 (insertion and deletion AUC), Fig. 6 (insertion and deletion curves for MNIST), and Fig. 7 (insertion and deletion curves for CIFAR-10).
@@
We compute this curve for each test point and average across the curves (as reported in Fig. 6 and Fig. 7). We limit the computation to the top 100 features for both datasets and report the values in percentages. Note that for MNIST, this is a complete ordering of all features (64), whereas for CIFAR-10, this is only a fraction of the 3072 dimensions. These are commonly used faithfulness measures from the literature: the higher the insertion AUC / the lower the deletion AUC, the higher the faithfulness of the model.
Moreover, for each test point, we compute the Spearman correlation rank from the scipy stats library, on the Shapley values, summed (in absolute value) across classes. We report the results in Table 7. This is a measure of agreement between the true Shapley values and the estimated Shapley values. The higher the Spearman rank correlation, the better the faithfulness of the approximation. Note, however, that due to the presence of many small values (near zero Shapley values) in image classification tasks, this measure may overemphasize incongruence between features.
@@
$\ell_{2}$ squared 0.0039 0.00222
$\ell_{2}$ modified 0.00427 0.00193
Communities Kernel kernel 0.00104
Matrix-Vec kernel 0.0214 0.0175
$\ell_{2}$ squared 0.0117 0.0118
$\ell_{2}$ modified 0.0131 0.0133
Least Squares kernel 1.3e-05 1.07e-05
$\ell_{2}$ squared 7.79e-06 7.21e-06
$\ell_{2}$ modified 8.67e-06 7.8e-06
Correlated Kernel kernel 0.00217
Matrix-Vec kernel 0.0155 0.00999
$\ell_{2}$ squared 0.0083 0.00782
$\ell_{2}$ modified 0.00988 0.00799
Least Squares kernel 4.89e-05 0.00105
$\ell_{2}$ squared 2.84e-05 0.000768
$\ell_{2}$ modified 3.19e-05 0.000817
@@
$\ell_{2}$ squared 0.00155 0.00889
$\ell_{2}$ modified 0.0016 0.00983
Independent Kernel kernel 0.000497
Matrix-Vec kernel 0.0135 0.00915
$\ell_{2}$ squared 0.00798 0.00671
$\ell_{2}$ modified 0.00944 0.0072
Least Squares kernel 1.11e-05 0.000231
$\ell_{2}$ squared 6.76e-06 0.00017
$\ell_{2}$ modified 7.73e-06 0.000191
@@
$\ell_{2}$ squared 3.28e-05 1.45e-05
$\ell_{2}$ modified 3.28e-05 1.54e-05
NHANES Kernel kernel 0.00253
Matrix-Vec kernel 0.00681 0.00281
$\ell_{2}$ squared 0.00588 0.00325
$\ell_{2}$ modified 0.00554 0.00287
Least Squares kernel 6.08e-05 0.00238
$\ell_{2}$ squared 5.02e-05 0.00221
$\ell_{2}$ modified 5.28e-05 0.00226
Table 2: Mean Squared Error for Different Sampling and Approximation Methods Across Various Datasets
@@
$\ell_{2}$ squared 0.00281 0.0015
$\ell_{2}$ modified 0.00245 0.00136
Communities Kernel kernel 0.000989
Matrix-Vec kernel 0.0195 0.0161
$\ell_{2}$ squared 0.0107 0.0105
$\ell_{2}$ modified 0.0121 0.0114
Least Squares kernel 1.19e-05 9.84e-06
$\ell_{2}$ squared 6.95e-06 6.48e-06
$\ell_{2}$ modified 7.98e-06 7.22e-06
Correlated Kernel kernel 0.00189
Matrix-Vec kernel 0.013 0.00887
$\ell_{2}$ squared 0.00755 0.00666
$\ell_{2}$ modified 0.00847 0.00683
Least Squares kernel 4.35e-05 0.000919
$\ell_{2}$ squared 2.55e-05 0.000653
$\ell_{2}$ modified 2.84e-05 0.000734
@@
$\ell_{2}$ squared 0.000874 0.00667
$\ell_{2}$ modified 0.00126 0.00799
Independent Kernel kernel 0.000452
Matrix-Vec kernel 0.0119 0.00805
$\ell_{2}$ squared 0.00678 0.00577
$\ell_{2}$ modified 0.00824 0.00642
Least Squares kernel 9.83e-06 0.000204
$\ell_{2}$ squared 6.13e-06 0.000148
$\ell_{2}$ modified 6.8e-06 0.000163
@@
$\ell_{2}$ squared 1.15e-05 1.25e-05
$\ell_{2}$ modified 8.11e-06 1.44e-05
NHANES Kernel kernel 0.00208
Matrix-Vec kernel 0.00512 0.00226
$\ell_{2}$ squared 0.00455 0.00258
$\ell_{2}$ modified 0.00456 0.0024
Least Squares kernel 4.86e-05 0.00197
$\ell_{2}$ squared 3.96e-05 0.00184
$\ell_{2}$ modified 4.06e-05 0.00191
@@
$\ell_{2}$ squared 0.00542 0.00317
$\ell_{2}$ modified 0.00586 0.00331
Communities Kernel kernel 0.00117
Matrix-Vec kernel 0.0238 0.02
$\ell_{2}$ squared 0.0129 0.013
$\ell_{2}$ modified 0.0146 0.0145
Least Squares kernel 1.42e-05 1.17e-05
$\ell_{2}$ squared 8.43e-06 7.82e-06
$\ell_{2}$ modified 9.37e-06 8.75e-06
Correlated Kernel kernel 0.00246
Matrix-Vec kernel 0.0177 0.0117
$\ell_{2}$ squared 0.00961 0.00901
$\ell_{2}$ modified 0.0112 0.00934
Least Squares kernel 5.52e-05 0.00119
$\ell_{2}$ squared 3.22e-05 0.000892
$\ell_{2}$ modified 3.64e-05 0.000932
@@
$\ell_{2}$ squared 0.00237 0.0144
$\ell_{2}$ modified 0.0024 0.0141
Independent Kernel kernel 0.000552
Matrix-Vec kernel 0.0154 0.0108
$\ell_{2}$ squared 0.0092 0.00784
$\ell_{2}$ modified 0.0109 0.00846
Least Squares kernel 1.26e-05 0.000268
$\ell_{2}$ squared 7.49e-06 0.00019
$\ell_{2}$ modified 8.78e-06 0.000212
@@
$\ell_{2}$ squared 3.28e-05 0.000131
$\ell_{2}$ modified 3.28e-05 3.3e-05
NHANES Kernel kernel 0.00302
Matrix-Vec kernel 0.00849 0.00343
$\ell_{2}$ squared 0.00754 0.00385
$\ell_{2}$ modified 0.00755 0.00368
Least Squares kernel 7.87e-05 0.00292
$\ell_{2}$ squared 6.01e-05 0.00281
$\ell_{2}$ modified 6.26e-05 0.00287
@@
$\ell_{2}$ squared 0.095 0.129 0.241
$\ell_{2}$ modified 0.098 0.15 0.233
Communities 16384 kernel 0.0207 0.0226 0.025
$\ell_{2}$ squared 0.0195 0.0213 0.0237
$\ell_{2}$ modified 0.0173 0.0193 0.0214
Correlated 16384 kernel 0.0153 0.0174 0.0196
$\ell_{2}$ squared 0.0173 0.0195 0.0227
$\ell_{2}$ modified 0.0141 0.0169 0.019
@@
$\ell_{2}$ squared 0.328 0.432 0.562
$\ell_{2}$ modified 0.29 0.393 0.577
Independent 16384 kernel 0.0166 0.0187 0.0205
$\ell_{2}$ squared 0.0203 0.0227 0.0252
$\ell_{2}$ modified 0.016 0.0181 0.0204

### v2

@@ Appendix F Methodology
In this section, we describe our estimators algorithmically. The unified theoretical framework can directly be implemented into an algorithmic framework, which we depict in Fig.˜5. The general procedure to generate the Shapley values in our framework requires three choices: (1) a sampling distributions on the index-sizes, (2) a strategy for sampling (with replacement, without replacement) and (3) an approximation method (least squares or matrix-vector). Least squares and matrix-vector estimation are reported in Section˜2. The missing detail is how the sampling procedure is implemented (this is the middle column in Fig.˜5). We report this in Algorithm˜1 for with replacement sampling and in Algorithm˜2.
[…]
Figure 5: The unified framework for estimating Shapley values with the proposed class of estimators. First, we define a distribution to apply to each bucket (i.e., to the selection of the bit vector to select - $p_{i}$ is the probability of sampling an item from bucket/coalition of size (or bit vector with Hamming weight) $i\in[d]$. Then we select a sampling strategy (with or without replacement). Finally, we select the estimation strategy. If we limit ourselves to $\ell_{2}$-squared and modified, and kernel distribution, this provides a total of $3\times 2\times 2=12$ estimators.
@@ F.1 With Replacement Estimators
Sampling with replacement to generate the sketch is a computationally efficient procedure that performs well in practice. However, if the number of samples $m>2^{d}$, the estimator will fail to compute exact Shapley values in general. We report the sampling procedure as implemented in our experimental evaluations in Algorithm˜1.
@@
1:$d$: number of features $m$: number of samples expected by user, $p$ choice of distribution on d1d-1 buckets, maxval maximum value before Poisson approximation is applied (our algorithm uses $10^{10})$ .
[…]
3:Redefine $p_{i}\leftarrow 2p_{i}$ for $i=1,\dotsc,\lfloor\frac{d-1}{2}\rfloor$.
4:First sample with replacement $\lfloor\frac{m}{2}\rfloor$ from indices $i=1,...,\lceil\frac{d-1}{2}\rceil$, $j\in[m]$ each with probability $p_{1},...,p_{\lceil\frac{d-1}{2}\rceil}$. Denote $m_{i}$ to be the number of times we observe the index (bucket) $i$ during sampling.
5:For each $i=1,\dotsc,\lceil\frac{d-1}{2}\rceil$, construct $\bm{b}_{j}\in\{0,1\}^{d}$ by sampling with replacement $m_{i}$ bitstrings of size $i$, which is equivalent to sampling without replacement $i$ indices from $[d]$, uniformly; and generate, the complement $\bar{\bm{b}}_{i,j}=1-\bm{b}_{i,j}$.
6:Construct $\bm{w}_{j}=\left((d-1)p_{j}\right)^{-1}$ and return:
@@ F.2 Without Replacement Estimators
At first glance, sampling based on the without replacement sampling scheme described in Section˜2.3 (see also Section˜A.5) requires flipping $2^{[d]}-2$ coins. However, for the sampling distributions described in Section˜2.2, given the size of a subset, the probability of picking any given subset of that size is constant. This observation is used in [MW25] to avoid flipping exponentially many coins. In particular, one can determine which coins are heads by first determining how many heads there will be of a given subset size, and then picking the resulting subsets of this size uniformly at random. We describe a variant of [MW25, Algorithm 2] in Algorithm˜2.
Algorithm 2 Sampling without Replacement (paired sampling, modified from [MW25])
1:$d$: number of features $m$: number of samples expected by user, $p$ choice of distribution on $d$ buckets, maxval maximum value before Poisson approximation is applied (our algorithm uses $10^{10})$ .
[…]
3:Redefine $p_{i}\leftarrow 2p_{i}$ for $i=1,\dotsc,\lfloor\frac{d-1}{2}\rfloor$.
4:Choose $\alpha$ such that
$\lfloor\frac{m}{2}\rfloor=\sum_{i=1}^{\lceil\frac{d-1}{2}\rceil}\min\left({d\choose i},\alpha p_{i}\right)$
@@
5:if ${d\choose i}<\texttt{maxval}$ then let
$m_{i}\leftarrow\text{Binomial}\left({d\choose i},\min(1,\alpha p_{i})\right)$
6:else let
$m_{i}\leftarrow\text{Poisson}(\alpha p_{i}),$
and let $\tilde{m}=\sum_{i=1}^{\lceil\frac{d-1}{2}\rceil}m_{i}$.
7:end if
8:Construct $j\in[m_{i}]$ bitstring arrays of $\bm{b}_{j}\in\{0,1\}^{d}$ of size $i\in[\lceil\frac{d-1}{2}\rceil]$ without replacement (e.g. using Fisher Yates shuffling or Algorithms 2,3 in [MW25]). If there is a middle bucket (i.e., $d$ is odd), fix $\bm{b}_{\lceil\frac{d-1}{2}\rceil,j}=1$ and sample without replacement from the remaining bitstrings. Then, generate the complement $\bar{\bm{b}}_{i,j}=1-\bm{b}_{i,j}$.
9:Construct $\bm{w}_{j}=\left((d-1)\min({d\choose j},\alpha p_{j})\right)^{-1}$ and return:
@@ Appendix G Experimental Details
We use publicly available datasets for reproducibility; choosing particularly those available through the shap for their popularity, ease of use and for a direct comparison with [MW25].
@@ G.1 Training Details
In this subsection, we detail the experimental design choices and hyperparameter for low Section˜3.1 and Section˜3.2, including implementation details, to promote reprodicibility.
@@ G.1.1 Low Dimensional Experiments
We refer to low dimensional experiments to the content of Section˜3.1. For each dataset we train a decision tree from the xgboost. Specifically, we use the XGBRegressor class with 100 estimators and maximum depth of 10 for each task. We replacing missing values with the mean for that feature in the dataset. Note that the goal is not to achieve competitive performance but to rapidly train a model where the Shapley values can be computed exactly and efficiently. The train test splits are ordered 80/20 splits for all datasets; we pick as query and baseline points the first data points of the test and train datasets respectively.
@@ G.1.2 High Dimensional Experiments
We refer to high dimensional experiments to the content of Section˜3.2. For the two classification tasks, we train a RandomForestClassifier from the sci-kit learn library. The random forest has maximum depth 15 and random state 42 for both tasks (MNIST and CIFAR-10). For both datasets, we pick as query and baseline points the first data points of the test and train datasets respectively.
For MNIST we use train test splits (80/20) with random state 42 using the train_test_split method on sci-kit learn. We achieve a test accuracy of $96.3\%$.
For the CIFAR-10 dataset, we use standard train test split from the torchvision library. We preprocess the input data by normalising. We achieve low accuracy $44.98\%$, as should be expected using basic tree-based models for CIFAR-10; however, using said models enables the exact computation of Shapley values, which we consider a more important aspect for the paper. We used paired, with replacement estimators with modified $\ell_{2}$ and $\bm{b}_{\alpha}$
@@ G.2 Datasets
Adult. Demographic information about individuals collected from the 1994 U.S. Census database. It is used to predict whether a person earns more than $\mathdollar 50,000$ dollars per year based on individual attributes: age, work, class, education, etc.,.
@@ G.2.2 Image Datasets
MNIST The MNIST dataset is a collection of handwritten digets (0-9); the classic task is to classify into their respective value. The dataset has $28\times 28(=748)$ dimensions. This is an incredibly popular dataset used for training and testing image classification algorithms.
CIFAR-10 CIFAR-10 dataset is a classification task dataset where a collection of 60,000 images, with $32\times 32\times 3(=3072)$ dimensions, are mapped to target clas: cars, airplanes, birds, cats, deer, dogs, frogs, horses, ships and trucks. The standard split for this dataset is 10,000 test images and 50,000 training images, which we use in our experiments. This is considered a relatively challenging dataset for boosted trees and feedforward neural networks; good performance is achieved, however, for convolutional neural networks.
@@ Appendix H Extended Experimental Results
The goal of this section is to report the numerical results from our experiments. We first report the extended experiments from Section˜3.1, followed by experiments in Section˜3.2. Importantly, we share tables and plots containing our results.
@@ H.1 Low Dimensional Experiments
In low dimensional experiments, as described in Appendix˜G and Section˜3.1, we compute the mean squared error (normalized by norm of exact Shapley values) as the median of 100 random seeds (0-99). We also report the average and interquartile ranges for each of the experiments. These results are summarized in Table˜2 (median), Table˜3 (lower quantile), Table˜4 (upper quantile), and Table˜5 ($b_{0}$ values). In each of those tables we report the values for selected number of samples: for IRIS we report $m=10$; Adult, California and Diabetes we show results for $m=64$; for Communities, Correlated, Independent, and NHANES we report $m=50000$.
@@ H.2 High Dimensional Experiments
For high dimensional experiments, as described in Appendix˜G and Section˜3.2, we compute the the mean squared error (normalized by norm of exact Shapley values) as the average of 10 test points from the respective datasets. We also report quantiles and median in Table˜6.
@@ H.3 Faithfulness Experiments
For each experiment in the high dimensional setting, we compute the insertion and deletion curves as reported in Table˜7 (insertion and deletion AUC), Fig.˜6 (insertion and deletion curves for MNIST), and Fig.˜7 (insertion and deletion curves for CIFAR-10).
@@
We compute this curve for each test point and average across the curves (as reported in Fig.˜6 and Fig.˜7). We limit the computation to the top 100 features for both datasets and report the values in percentages. Note that for MNIST, this is a complete ordering of all features (64), whereas for CIFAR-10, this is only a fraction of the 3072 dimensions. These are commonly used faithfulness measures from the literature: the higher the insertion AUC / the lower the deletion AUC, the higher the faithfulness of the model.
Moreover, for each test point, we compute the Spearman correlation rank from the scipy stats library, on the Shapley values, summed (in absolute value) across classes. We report the results in Table˜7. This is a measure of agreement between the true Shapley values and the estimated Shapley values. The higher the Spearman rank correlation, the better the faithfulness of the approximation. Note, however, that due to the presence of many small values (near zero Shapley values) in image classification tasks, this measure may overemphasize incongruence between features.
@@
$\ell_{2}$-squared 0.0039 0.00222
modified $\ell_{2}$ 0.00427 0.00193
Communities Kernel kernel 0.000810
Matrix-Vec kernel 0.009871 0.007950
$\ell_{2}$-squared 0.009174 0.026858
modified $\ell_{2}$ 0.007645 0.018672
Least Squares kernel 0.000887 0.000719
$\ell_{2}$-squared 0.000479 0.000301
modified $\ell_{2}$ 0.000554 0.000315
Correlated Kernel kernel 0.000620
Matrix-Vec kernel 0.007933 0.005208
$\ell_{2}$-squared 0.009155 0.002495
modified $\ell_{2}$ 0.006804 0.002772
Least Squares kernel 0.000252 0.000169
$\ell_{2}$-squared 0.000146 0.000064
modified $\ell_{2}$ 0.000172 0.000072
@@
$\ell_{2}$-squared 0.00155 0.00889
modified $\ell_{2}$ 0.0016 0.00983
Independent Kernel kernel 0.000480
Matrix-Vec kernel 0.006640 0.004847
$\ell_{2}$-squared 0.005631 0.004184
modified $\ell_{2}$ 0.005024 0.003773
Least Squares kernel 0.000673 0.000502
$\ell_{2}$-squared 0.000401 0.000370
modified $\ell_{2}$ 0.000487 0.000377
@@
$\ell_{2}$-squared 3.28e-05 1.45e-05
modified $\ell_{2}$ 3.28e-05 1.54e-05
NHANES Kernel kernel 0.000597
Matrix-Vec kernel 0.011273 0.008429
$\ell_{2}$-squared 0.013369 0.009053
$\ell_{2}$-modified 0.009980 0.007377
Least Squares kernel 0.002637 0.002153
$\ell_{2}$-squared 0.001505 0.00221
modified $\ell_{2}$ 0.001674 0.000428
Table 2: Mean Squared Error for Different Sampling and Approximation Methods Across Various Datasets (best relative MSE performance marked in bold).
@@
$\ell_{2}$-squared 0.00281 0.0015
modified $\ell_{2}$ 0.00245 0.00136
Communities Kernel kernel 0.000735
Matrix-Vec kernel 0.009144 0.007282
$\ell_{2}$-squared 0.008249 0.025044
modified $\ell_{2}$ 0.006825 0.016340
Least Squares kernel 0.000785 0.000659
$\ell_{2}$-squared 0.000436 0.000276
modified $\ell_{2}$ 0.000504 0.000284
Correlated Kernel kernel 0.000547
Matrix-Vec kernel 0.006926 0.004395
$\ell_{2}$-squared 0.008373 0.002188
modified $\ell_{2}$ 0.006078 0.002340
Least Squares kernel 0.000219 0.000149
$\ell_{2}$-squared 0.000129 0.000058
modified $\ell_{2}$ 0.000149 0.000062
@@
$\ell_{2}$-squared 0.000874 0.00667
modified $\ell_{2}$ 0.00126 0.00799
Independent Kernel kernel 0.000435
Matrix-Vec kernel 0.005925 0.004309
$\ell_{2}$-squared 0.004942 0.003723
modified $\ell_{2}$ 0.004517 0.003284
Least Squares kernel 0.000581 0.000431
$\ell_{2}$-squared 0.000353 0.000322
modified $\ell_{2}$ 0.000423 0.000333
@@
$\ell_{2}$-squared 1.15e-05 1.25e-05
modified $\ell_{2}$ 8.11e-06 1.44e-05
NHANES Kernel kernel 0.000505
Matrix-Vec kernel 0.010318 0.007504
$\ell_{2}$-squared 0.011346 0.008037
modified $\ell_{2}$ 0.008947 0.006688
Least Squares kernel 0.002385 0.001926
$\ell_{2}$-squared 0.001368 0.00221
modified $\ell_{2}$ 0.001540 0.000375
@@
$\ell_{2}$-squared 0.00542 0.00317
modified $\ell_{2}$ 0.00586 0.00331
Communities Kernel kernel 0.000877
Matrix-Vec kernel 0.010869 0.008674
$\ell_{2}$-squared 0.009949 0.029046
modified $\ell_{2}$ 0.008383 0.019998
Least Squares kernel 0.000968 0.000787
$\ell_{2}$-squared 0.000522 0.000335
modified $\ell_{2}$ 0.000597 0.000346
Correlated Kernel kernel 0.000721
Matrix-Vec kernel 0.009207 0.005980
$\ell_{2}$-squared 0.010407 0.002896
modified $\ell_{2}$ 0.007889 0.003279
Least Squares kernel 0.000294 0.000195
$\ell_{2}$-squared 0.000159 0.000073
modified $\ell_{2}$ 0.000184 0.000081
@@
$\ell_{2}$-squared 0.00237 0.0144
modified $\ell_{2}$ 0.0024 0.0141
Independent Kernel kernel 0.000551
Matrix-Vec kernel 0.007710 0.005508
$\ell_{2}$-squared 0.006441 0.004724
modified $\ell_{2}$ 0.005656 0.004196
Least Squares kernel 0.000745 0.000558
$\ell_{2}$-squared 0.000463 0.000413
modified $\ell_{2}$ 0.000529 0.000444
@@
$\ell_{2}$-squared 3.28e-05 0.000131
modified $\ell_{2}$ 3.28e-05 3.3e-05
NHANES Kernel kernel 0.000704
Matrix-Vec kernel 0.012395 0.009490
$\ell_{2}$-squared 0.014504 0.009863
modified $\ell_{2}$ 0.010979 0.008374
Least Squares kernel 0.002917 0.002366
$\ell_{2}$-squared 0.001668 0.00221
modified $\ell_{2}$ 0.001888 0.000471
@@
$\ell_{2}$-squared 0.095 0.129 0.241
modified $\ell_{2}$ 0.098 0.15 0.233
Communities 50000 kernel 0.000765 0.000851 0.000920
$\ell_{2}$-squared 0.000555 0.000600 0.000672
modified $\ell_{2}$ 0.000595 0.000637 0.000695
Correlated 50000 kernel 0.000495 0.000566 0.000644
$\ell_{2}$-squared 0.000402 0.000448 0.000528
modified $\ell_{2}$ 0.000431 0.000493 0.000562
@@
$\ell_{2}$-squared 0.328 0.432 0.562
modified $\ell_{2}$ 0.29 0.393 0.577
Independent 50000 kernel 0.000435 0.000510 0.000541
$\ell_{2}$-squared 0.000368 0.000426 0.000487
modified $\ell_{2}$ 0.000407 0.000446 0.000494